# The Unpaid Toll: Quantifying the Public Health Impact of AI

### v1

@@ 4 Results
We now present our estimates of the public health impacts caused by the U.S. data centers in aggregate and by training a large generative AI model at specific locations. Our results demonstrate that in 2030, the scope-2 pollutants of U.S. data centers alone could cause, among others, approximately 600,000 asthma symptom cases and 1,300 premature deaths, exceeding 1/3 of asthma deaths in the U.S. each year [47]. The overall public health costs of U.S. data centers could rival or even exceed those of on-road emissions of the largest U.S. states such as California. Moreover, depending on the locations, training an AI model of the Llama-3.1 scale can produce an amount of air pollutants equivalent to driving a passenger car for more than 10,000 LA-NYC round trips, resulting in a health cost that even exceeds 120% of the training electricity cost. Importantly, the health costs are disproportionately distributed across counties and communities, particularly affecting low-income counties that could experience more than 200x per-household health costs than others.
4.1 Public Health Impact of U.S. Data Centers in 2023
We first show in Table 1 the public health cost of U.S. data centers in 2023 as a reference. [^5] We use the “mid (low, high)” format to represent the midrange, low and high estimates offered by COBRA. When presenting a single value or a ratio (e.g., health-to-electricity cost ratio), we use the midrange by default. Even at the beginning of the generative AI boom, the U.S. data centers have already resulted in a total public health cost of about \$5.6 billion, or \$39.7 per household, in 2023. This is equivalent to 43% of the data centers’ total electricity cost. By considering marginal attribution, the U.S. data centers’ public health cost increases to about \$7.6 billion in 2023, due to the heavy reliance on fossil fuels by many marginal generators [70]. This suggests that, by powering the U.S. data centers using alternative energy sources (e.g., geothermal) off the main grid, the U.S. could have seen a public health benefit of \$7.6 billion in 2023. Additional results can be found in Appendix B, including county-wide total and per-household health costs that demonstrate the uneven distribution of health impacts across different communities.
Table 1: The public health cost of U.S. data centers in 2023.
Attribution Method Electricity (TWh) Electricity Cost (billion \$) Mortality Health Cost (billion \$) % of Electricity Cost Per-Household Health Cost (\$) Months of Electricity Bill % of CA On-road Health Cost
Average 152.1 13.0 360 (270, 460) 5.6 (4.2, 7.0) 43% 39.7 (29.6, 49.8) 0.29 (0.22, 0.36) 35%
Marginal 152.1 13.0 490 (360, 620) 7.6 (5.7, 9.4) 58% 53.8 (40.5, 67.2) 0.39 (0.30, 0.49) 47%
@@
Mobile sources, including vehicles, marine engines, and generators, collectively account for more than half of the air pollutants in the U.S., with vehicles being a primary contributor [80, 81]. Thus, we contextualize the data centers’ public health cost by comparing it to that of on-road emissions of California, which has about 35 million registered vehicles and exhibits the highest public health cost resulting from on-road emissions among all the U.S. states [49, 46]. On-road emissions are categorized as the “Highway Vehicles” sector in COBRA and include both tailpipe exhaust and tire and brake wear. The details of calculating on-road emissions and the corresponding health costs are available in Appendix A.1. We see from Table 1 that in 2023, the total public health cost of U.S. data centers exceeds 1/3 of that of California’s on-road emissions.
4.2 Public Health Impact of U.S. Data Centers in 2030
This section presents our projections of the public health cost of the U.S. data centers in 2030.
We first show in Fig. 2 the health costs of U.S. data centers and compare them with top-3 state on-road emissions in 2030 by using different extrapolation methods. More detailed results are available in Table 2. Due to the tightening air pollutant regulations [82], the health costs of on-road emissions — a primary source of air pollutants in the U.S. — have generally decreased from 2016 to 2030. In contrast, the surging demand for AI data centers in the U.S. has outweighed the power plant emission efficiency improvement, potentially quadrupling the public health cost from 2023 to 2030. Under McKinsey’s projection with a medium growth rate, the scope-2 pollutants of U.S. data centers in 2030 alone could cause, among others, approximately 600,000 asthma symptom cases and 1,300 deaths, exceeding 1/3 of asthma deaths in the U.S. each year [47]. Importantly, the public health costs of U.S. data centers could rival or even exceed those of on-road emissions of the largest U.S. states including California, suggesting a need for urgent attention to the health impact of U.S. data centers beyond on-road emissions.
(a) Linear extrapolation
(b) Exponential extrapolation
(c) 2028-to-2030 unchanged
Figure 2: The health costs of U.S. data centers and top-3 state on-road emissions from 2016 to 2030 based on different extrapolations for 2030 baseline emissions.
Table 2: The public health cost of U.S. data centers in 2030 based on EPRI’s energy demand projection [5]. “” denotes McKinsey’s projection under a medium growth rate [4].
Extrapolation Method Projected Growth Electricity (TWh) Electricity Cost (billion \$) Mortality Health Cost (billion \$) % of Electricity Cost Per-Household Health Cost (\$) Months of Electricity Bill % of CA On-road Health Cost
Unchanged Low 196.3 16.8 490 (360, 610) 8.3 (6.3, 10.3) 49% 55.3 (41.9, 68.8) 0.40 (0.31, 0.50) 43%
Moderate 214.0 18.3 530 (390, 660) 9.0 (6.8, 11.1) 49% 59.9 (45.4, 74.5) 0.44 (0.33, 0.54) 47%
High 296.4 25.4 710 (530, 890) 12.1 (9.2, 15.1) 48% 80.9 (61.3, 100.6) 0.59 (0.45, 0.74) 63%
Higher 403.9 34.6 940 (700, 1180) 16.0 (12.1, 19.8) 46% 106.7 (80.8, 132.6) 0.78 (0.59, 0.97) 84%
Medium 519.3 44.5 1210 (900, 1510) 20.5 (15.5, 25.5) 46% 137.2 (103.9, 170.5) 1.00 (0.76, 1.25) 107%
Linear Low 196.3 16.8 500 (370, 630) 8.5 (6.4, 10.5) 50% 56.5 (42.6, 70.5) 0.41 (0.31, 0.51) 61%
Moderate 214.0 18.3 540 (400, 680) 9.2 (6.9, 11.4) 50% 61.3 (46.2, 76.3) 0.45 (0.34, 0.56) 66%
High 296.4 25.4 730 (540, 920) 12.4 (9.3, 15.4) 49% 82.7 (62.4, 103.1) 0.60 (0.46, 0.75) 90%
Higher 403.9 34.6 960 (710, 1210) 16.3 (12.3, 20.3) 47% 109.0 (82.2, 135.9) 0.80 (0.60, 0.99) 118%
Medium 519.3 44.5 1230 (920, 1550) 21.0 (15.8, 26.1) 47% 140.2 (105.7, 174.7) 1.02 (0.77, 1.28) 152%
Exponential Low 196.3 16.8 510 (380, 640) 8.7 (6.6, 10.8) 52% 58.0 (43.8, 72.3) 0.42 (0.32, 0.53) 53%
Moderate 214.0 18.3 550 (410, 700) 9.4 (7.1, 11.7) 51% 62.9 (47.4, 78.3) 0.46 (0.35, 0.57) 57%
High 296.4 25.4 750 (560, 940) 12.7 (9.6, 15.8) 50% 84.9 (64.1, 105.7) 0.62 (0.47, 0.77) 78%
Higher 403.9 34.6 990 (730, 1240) 16.7 (12.6, 20.9) 48% 111.9 (84.4, 139.4) 0.82 (0.62, 1.02) 102%
Medium 519.3 44.5 1270 (940, 1590) 21.5 (16.2, 26.8) 48% 143.9 (108.6, 179.2) 1.05 (0.79, 1.31) 132%
(a) Per-household health cost map
(b) CDF of county-level per-household health cost
State County Per-household Health Cost (\$) Months of Electricity Bills County-to-nation Income Ratio WV Mason 986.4 (782.3, 1190.5) 7.1 (5.6, 8.6) 0.71 OH Meigs 981.1 (744.5, 1217.7) 7.9 (6.0, 9.8) 0.62 WV Marion 977.0 (783.2, 1170.9) 7.1 (5.7, 8.5) 0.80 WV Marshall 974.0 (759.0, 1189.0) 7.0 (5.5, 8.6) 0.77 OH Gallia 939.6 (710.7, 1168.6) 7.5 (5.7, 9.4) 0.74 WV Taylor 890.1 (722.9, 1057.2) 6.4 (5.2, 7.6) 0.70 PA Fayette 874.9 (694.0, 1055.7) 6.1 (4.8, 7.4) 0.74 WV Brooke 861.6 (660.5, 1062.7) 6.2 (4.8, 7.7) 0.69 PA Greene 837.9 (687.2, 988.5) 5.9 (4.8, 6.9) 0.88 WV Jackson 753.2 (610.5, 895.9) 5.4 (4.4, 6.5) 0.73
(c) Top-10 counties by per-household health cost
County Per-household Health Cost (\$) Months of Electricity Bills County-to-nation Income Ratio Emporia City 510.4 (369.2, 651.6) 3.6 (2.6, 4.6) 0.55 Lancaster 509.5 (405.5, 613.6) 3.6 (2.9, 4.3) 0.83 Staunton City 502.7 (386.3, 619.2) 3.5 (2.7, 4.4) 0.79 Buena Vista City 490.5 (378.8, 602.3) 3.5 (2.7, 4.3) 0.65 Highland 466.0 (367.9, 564.1) 3.3 (2.6, 4.0) 0.76 Buchanan 463.2 (362.3, 564.0) 3.3 (2.6, 4.0) 0.53 Fairfax City 462.8 (344.9, 580.8) 3.3 (2.4, 4.1) 1.71 Dickenson 454.0 (354.2, 553.9) 3.2 (2.5, 3.9) 0.53 Tazewell 441.8 (344.8, 538.8) 3.1 (2.4, 3.8) 0.62 Colonial Heights City 435.5 (340.3, 530.7) 3.1 (2.4, 3.7) 0.96
(d) Top-10 counties in Virginia by per-household health cost
Figure 3: The county-level per-household health cost of U.S. data centers in 2030 based on exponential extrapolation of baseline emissions (McKinsey’s medium-growth forecast). The income data is based on the 2018-2022 American Community Survey 5-year estimates provided by [83].
Next, we show in Fig. 3 the county-level per-household health cost of U.S. data centers in 2030 based on exponential extrapolation under McKinsey’s medium-growth forecast. We see that the health cost is highly disproportionately distributed across different counties and communities, particularly affecting low-income communities. The ratio of the highest county-level per-household health cost to the lowest cost could be more than 200. Crucially, all the top-10 counties in the U.S. and 9 out of top-10 counties in Virginia (which has the largest concentration of data centers in the U.S. [5, 4]) have lower median household incomes than the national median value. Moreover, many of the hardest-hit communities do not have large data centers or directly receive economic benefits from AI data centers such as tax revenues. Yet, compared to the national average of about 1 month of electricity bill per year, the households in these communities could each suffer from health impacts equivalent to up to $\sim$8 months of their electricity bills. The high degree of disparity across different communities in terms of the public health cost suggests that we must examine the local and regional health impacts of AI data centers and improve public health equity to enable truly responsible AI.
(a) Health cost map
(b) CDF of county-level health cost
State County Health Cost (million \$) IL Cook 309.6 (214.4, 404.7) PA Allegheny 266.4 (204.7, 328.1) TX Harris 262.0 (182.9, 341.1) MI Wayne 171.8 (123.0, 220.6) PA Philadelphia 169.4 (118.0, 220.8) OH Hamilton 166.5 (127.3, 205.7) NY Kings 165.2 (105.2, 225.2) OH Franklin 152.0 (112.2, 191.8) OH Cuyahoga 151.1 (110.6, 191.5) NY Queens 138.6 (90.7, 186.5)
(c) Top-10 counties by total health cost
Figure 4: The county-level total health cost of U.S. data centers in 2030 based on exponential extrapolation of baseline emissions (McKinsey’s medium-growth forecast).
We also show the county-level total public health cost in Fig. 4. Compared to the per-household health cost distribution in Fig. 3, the county-level total health cost distribution is more aligned with the population distribution — despite the low per-household health cost, populous counties in California have a high total health cost. Nonetheless, some less populous counties (e.g., Hamilton County, Ohio) near coal and/or natural gas power plants are still significantly impacted and even more so than those (e.g., Loudoun County, Virginia) that have high concentrations of data centers.
@@
Table 3: The public health cost of training an AI model of the Llama-3.1 scale in Meta’s U.S. data centers.
@@
Location Electricity Price (¢/kWh) Electricity Cost (million \$) Health Cost (million \$) % of Electricity Cost Emission (Metric Ton)
@@ 4.3 Public Health Impact of Generative AI Training
We now study the health impact of training a generative AI model. Specifically, we consider the training of an LLM and assume that the electricity consumption is the same as training Llama-3.1 recently released by Meta [84]. While we use Meta’s Llama-3.1 training electricity consumption and U.S. data center locations as an example, our results should be interpreted as the estimated public health impact of training a general LLM with a comparable scale of Llama-3.1.
@@
5 Our Recommendations
We provide our recommendations to address the increasing public health impact of AI.
Recommendation 1: Standardization of Reporting Protocols
Despite their immediate and tangible impacts on public health, criteria air pollutants have been entirely overlooked in AI model cards and sustainability reports published by technology companies [45, 44, 10]. The absence of such critical information adds substantial challenges to accurately identifying specific AI data centers as a key root cause of public health burdens and could potentially pose hidden risks to public health. To enhance transparency and lay the foundation for truly responsible AI, we recommend standardization of reporting protocols for criteria air pollutants and the public health impacts across different regions. Concretely, criteria air pollutants can be categorized into three different scopes (Section 2.2), and reported following the greenhouse gas protocol widely adopted by technology companies [45, 10, 85].
Recommendation 2: Health-informed AI

### v2

@@ 5 Results
We now present our estimates of the public health impacts caused by the U.S. data centers in aggregate and by training a large generative AI model at specific locations. We focus on the contiguous U.S., which hosts nearly half of the world’s data centers [38], and simply refer to it as the U.S. For consistency with COBRA, cities considered county-equivalents for census purposes are also referred to as “counties” in our paper. All our monetary values are for one year (or one computing task if applicable) and in 2023 U.S. dollars as used by COBRA.
@@
Our results demonstrate that in 2028, the total scope-1 and scope-2 pollutants of U.S. data centers alone could cause, among others, approximately 600,000 asthma symptom cases and 1,300 premature deaths, exceeding 1/3 of asthma deaths in the U.S. each year [40]. The overall public health costs could reach more than \$20 billion, rival or even top those of on-road emissions of the largest U.S. states such as California with $\sim$35 million registered vehicles [41]. Importantly, the health costs are unevenly distributed across counties and communities, particularly affecting low-income counties that could experience approximately 200x per-household health costs than others. Moreover, depending on the locations, training an AI model of the Llama-3.1 scale can produce an amount of air pollutants equivalent to driving a passenger car for more than 10,000 LA-NYC round trips, resulting in a health cost that even exceeds 120% of the training electricity cost.
5.1 Public Health Impact of U.S. Data Centers
We now present our public health impact analysis for U.S. data centers in aggregate, beginning with the historical analysis from 2019 to 2023 and highlighting the uneven distribution of health impacts, followed by a 2028 projection. The overall trend is shown in Fig. 4, which demonstrates a significant increase in the public health impact of U.S. data centers from 2023 to 2028. Specifically, the surging demand for AI data centers in the U.S. has outweighed the power plant emission efficiency improvement, potentially tripling the public health cost from 2023 to 2028.
We present the details in Table 1 and see that scope-2 health cost dominates the scope-1 cost. [^6] We use the “mid (low, high)” format to represent the midrange, low and high estimates offered by COBRA. When presenting a single value or a ratio (e.g., health-to-electricity cost ratio), we use the midrange by default. This suggests that while using alternative fuels for onsite generators can help improve local air quality around data centers, greater health benefits can be achieved by powering data centers with pollutant-free electricity sources.
@@
Mobile sources, including vehicles, marine engines, and generators, collectively account for more than half of the air pollutants in the U.S., with vehicles being a primary contributor [89, 90]. Thus, to contextualize the public health costs of data centers, we compare them to on-road emissions in the three largest U.S. states (California, Texas, and Florida). In particular, California, which has about 35 million registered vehicles, exhibits the highest public health cost from on-road emissions among all U.S. states [41, 39]. In the COBRA model, on-road emissions are categorized under the “Highway Vehicles” sector and include both tailpipe exhaust and tire and brake wear. The details of calculating on-road emissions and the corresponding health costs are provided in Appendix A.
As shown in Fig. 4, in 2023, the total public health cost of U.S. data centers is 42% of that from California’s on-road emissions. Due to the tightening air pollutant regulations [91], the health costs of on-road emissions have generally decreased from 2019 to 2028. However, with rapid growth, the public health impact of U.S. data centers is projected to rival or even surpass that of California’s on-road emissions by 2028, underscoring the need for greater attention to the growing public health impacts of the data center industry.
Figure 4: The public health costs of U.S. data centers and top-3 state on-road emissions from 2019 to 2023 and the 2028 projection based on the Lawrence Berkeley National Lab’s report [4]. The cost for U.S. data centers includes scope-1 and scope-2 impacts. The “High” and “Low” represent the high and low growth rates considered in [4].
@@
Table 1: The public health cost of U.S. data centers from 2019 to 2023 and projection in 2028
Year Electricity (TWh) Electricity Cost (billion \$) Scope Mortality Health Cost (billion \$) Per-Household Health Cost (\$) % of CA On-road Health Cost
2019 90.28 7.73 Scope-1 7 (6, 8) 0.11 (0.10, 0.13) 0.84 (0.72, 0.97) 1%
Scope-2 189 (138, 240) 2.92 (2.18, 3.66) 21.51 (16.05, 26.97) 14%
Total 196 (144, 248) 3.03 (2.27, 3.79) 22.36 (16.77, 27.94) 15%
2020 105.56 9.04 Scope-1 9 (8, 11) 0.15 (0.13, 0.17) 1.11 (0.94, 1.27) 1%
Scope-2 233 (171, 296) 3.60 (2.68, 4.51) 26.30 (19.62, 32.98) 18%
Total 243 (179, 307) 3.75 (2.81, 4.69) 27.41 (20.56, 34.25) 19%
2021 127.78 10.94 Scope-1 13 (11, 15) 0.20 (0.17, 0.24) 1.48 (1.26, 1.70) 1%
Scope-2 280 (205, 355) 4.31 (3.22, 5.41) 31.24 (23.31, 39.17) 23%
Total 293 (216, 370) 4.52 (3.39, 5.64) 32.72 (24.57, 40.88) 24%
2022 151.39 12.97 Scope-1 21 (17, 24) 0.33 (0.28, 0.38) 2.40 (2.04, 2.76) 2%
Scope-2 330 (242, 418) 5.08 (3.79, 6.37) 36.44 (27.18, 45.70) 29%
Total 351 (259, 443) 5.41 (4.07, 6.75) 38.84 (29.22, 48.46) 31%
2023 176.39 15.11 Scope-1 32 (26, 37) 0.51 (0.43, 0.59) 3.65 (3.08, 4.21) 3%
Scope-2 401 (294, 508) 6.16 (4.59, 7.73) 43.83 (32.69, 54.97) 39%
Total 433 (320, 546) 6.67 (5.03, 8.32) 47.48 (35.77, 59.19) 42%
2019 651.40 55.79 Scope-1 82 (68, 95) 1.32 (1.12, 1.52) 9.49 (8.05, 10.92) 1%
to Scope-2 1434 (1050, 1818) 22.07 (16.46, 27.67) 159.32 (118.85, 199.80) 24%
2023 Total 1516 (1118, 1913) 23.38 (17.58, 29.19) 168.81 (126.90, 210.72) 25%
2028 (Low) 325.00 27.84 Scope-1 54 (46, 63) 0.90 (0.77, 1.03) 6.11 (5.22, 7.00) 5%
Scope-2 650 (483, 818) 10.78 (8.14, 13.41) 73.29 (55.37, 91.21) 61%
Total 705 (529, 880) 11.67 (8.91, 14.44) 79.40 (60.59, 98.21) 67%
2028 (High) 580.00 49.68 Scope-1 97 (82, 112) 1.61 (1.37, 1.84) 10.94 (9.35, 12.53) 9%
Scope-2 1165 (865, 1464) 19.29 (14.58, 24.01) 131.23 (99.14, 163.31) 110%
Total 1262 (947, 1576) 20.90 (15.95, 25.85) 142.16 (108.49, 175.84) 119%
5.1.1 Historical analysis: 2019-2023.
Table 1 shows the public health cost of U.S. data centers from 2019 to 2023 as a reference. Even at the beginning of the generative AI boom, the U.S. data centers have already resulted in a total public health cost of about \$6.7 billion, or \$47.5 per household, in 2023. This is equivalent to approximately 44% of the data centers’ total electricity cost.
Next, we show in Fig. 5 the county-level total public health cost of U.S. data centers from 2019 to 2023, which exhibits significant spatial variability. In particular, populated counties located downwind of power plants supplying electricity to data centers tend to experience higher health costs, reflecting the transport of air pollutants across regions. The cumulative distribution function (CDF) in Fig. 5b highlights that while many counties incur relatively low health costs, a small fraction of counties bear substantially higher impacts. Table 5c further identifies the top-10 counties with the highest total health costs, illustrating how local population density and proximity to power generation infrastructure combine to amplify public health risks in specific communities.
State County Health Cost (million \$) IL Cook 358.5 (245.7, 471.3) PA Allegheny 322.4 (246.3, 398.6) TX Harris 306.7 (208.5, 404.8) OH Hamilton 195.5 (147.8, 243.1) PA Philadelphia 193.7 (137.5, 250.0) MI Wayne 188.6 (133.2, 244.1) OH Cuyahoga 174.7 (125.7, 223.8) TX Dallas 174.4 (131.2, 217.6) OH Franklin 166.2 (121.8, 210.6) NY Kings 164.5 (102.7, 226.3)
Figure 5: The county-level total health cost of U.S. data centers from 2019 to 2023. (a) Health cost map; (b) CDF of county-level health cost; (c) Top-10 counties by total health cost.
5.1.2 Uneven distribution of data centers’ public health impacts.
Next, Fig. 6 presents the county-level per-household total health costs attributable to U.S. data centers from 2019 to 2023. The results reveal a highly disproportionate distribution of health impacts across counties, with low-income communities particularly affected. The ratio of the highest to lowest county-level per-household health cost reaches approximately 200. Notably, all of the top 10 counties with the highest per-household costs have median household incomes below the national median. The high degree of disparity across different communities in terms of the public health cost suggests that we need to carefully examine the local and regional health impacts of data centers and improve public health equity to enable truly responsible computing.
Furthermore, many of the hardest-hit communities neither host large data centers nor directly benefit economically from AI data centers, such as through tax revenues. For example, several counties in West Virginia are among the most affected, because many coal-fired power plants in West Virginia are supplying electricity to data centers in the neighboring state of Virginia [5, 4]. By contrast, despite hosting a large number of data centers and incurring high total health costs due to its large population, California’s clean power grid results in some of the lowest per-household health impacts in the country.
State County Per-household Health Cost  (\$) IR WV Marion 1218.3 (978.0, 1458.5) 0.80 WV Mason 1139.0 (897.6, 1380.4) 0.71 OH Meigs 1123.1 (840.8, 1405.5) 0.62 OH Gallia 1107.6 (828.1, 1387.2) 0.74 WV Marshall 1083.8 (831.2, 1336.5) 0.77 WV Taylor 1052.5 (853.1, 1252.0) 0.70 PA Fayette 992.6 (782.0, 1203.3) 0.74 PA Greene 944.9 (770.3, 1119.5) 0.88 WV Brooke 918.8 (693.2, 1144.4) 0.69 WV Jackson 871.9 (703.8, 1039.9) 0.73
Figure 6: The county-level per-household health cost of U.S. data centers from 2019 to 2023. (a) Per-household health cost map; (b) CDF of county-level per-household health cost; (c) Top-10 counties by per-household health cost. IR represents “County-to-Nation Per-Household Median Income Ratio.”
5.1.3 Projection for 2028.
According to a recent Lawrence Berkeley National Laboratory (LBNL) report [4], the U.S. data center electricity consumption is expected to increase from 4.4% of the total national electricity use in 2023 to 6.7–12.0% in 2028, depending on the growth trajectory of AI adoption. At the same time, the projected rise in peak power demand is accompanied by massive installations of onsite backup diesel generators to ensure reliability during grid contingencies [60].
This substantial growth in electricity demand and onsite generation is expected to offset, and in fact outweigh, the gradual pollution emission intensity reductions anticipated from the power sector. As a result, the total public health costs attributable to data center operations are projected to potentially triple from 2023 to 2028. Quantitatively, based on the low- and high-growth scenarios considered in [4], the total public health impact of U.S. data centers is estimated to reach \$11.7 billion and \$20.9 billion in 2028, respectively. Under the high-growth scenario, the resulting health burden could rival or exceed that of on-road emissions in the largest U.S. state. This highlights the growing health externalities of U.S. data centers at a national scale.
Table 2: The public health cost of training a large AI model in selected U.S. data centers.
@@
Location Electricity Price (¢/kWh) Electricity (million \$) Health Cost (million \$) % of Electricity Cost Emission (Metric Ton)
@@ 5.2 Public Health Impact of Generative AI Training
We now study the health impact of a specific computing task. Specifically, we consider the training of an LLM and assume that the electricity consumption is the same as training Llama-3.1 recently released by Meta [92]. As the scope-2 impact is dominant and the power allocated to train Llama-3.1 is unknown to determine scope-1 impacts, we focus on scope-2 health costs associated with the electricity consumption. While we use Meta’s Llama-3.1 training electricity consumption and U.S. data center locations as an example, our results should be interpreted as the estimated public health impact of training a general LLM with a comparable scale of Llama-3.1.
@@
5.3 Location-Dependent Public Health Impacts of Two Technology Companies
We further highlight locational dependency of public health impacts by considering two major technology companies’ U.S. data center locations in 2023, excluding their leased colocation data centers whose locations are proprietary. We name these two companies A and B, respectively. These two companies do not have same data center locations. While company B discloses its per-location electricity usage [37], company A does not. Thus, we uniformly distribute company B’s North America electricity consumption over its U.S. data center locations based on its latest sustainability report [10]. We consider location-based emission accounting without taking into account renewable energy credits these two companies apply to offset their grid electricity consumption (see “Location-based emission” in Section 4.1.2).
We see from Fig. 7 that the two companies have significant differences in terms of the per-household health cost distribution and most-affected counties. This is primarily due to the two companies’ different data center locations, and highlights the locational dependency of public health impacts. That is, unlike carbon emissions that have a similar effect regardless of the emission source locations, the public health impact of criteria air pollutants heavily depends on the location of the emission source. Thus, technology companies should account for public health impacts when deciding where they build data centers, where they get electricity for their data centers, and where they install onsite renewables in order to best mitigate the adverse health effects.
(a) Per-household health cost (Company A)
(b) Per-household health cost (Company B)
Figure 7: The county-level per-household health cost of two U.S. technology companies in 2023.
6 Health-Informed Computing: Addressing Data Centers’ Public Health Impact
In this section, we present Health-Informed Computing, a framework that explicitly incorporates public health impacts as a key optimization objective and strategically manages data center workloads to minimize adverse health outcomes while supporting broader sustainability goals.
To mitigate the public health impact of computing, one straightforward approach is to focus solely on reducing the energy consumption (e.g., reducing AI model sizes [93, 94]). While reducing energy consumption is beneficial, overlooking the downstream public health impact of where and when energy is produced does not necessarily lead to minimized health burdens. For example, Table 2 demonstrates a 10x difference in health costs for training the same AI model across different locations. This highlights that health-informed and energy-aware computing, when combined, offer complementary benefits, leading to better public health outcomes.
6.1 Opportunities for Health-Informed Computing